casie runs two and a half miles every day. after running seven eighths of a mile, how far does he have left to run? What is this question asking me to do
Answer by stanbon(75887)

casie runs two and a half miles every day. after running seven eighths of a mile, how far does he have left to run? What is this question asking me to do
----
total distance - distance already run = distance to be still run
------
2.5 - (7/8) = d
------
Multiply thru by 8 to get:
20 - 7 = 8d
---
distance to still be run = 13/8 mile
----------------
Cheers,
Stan H.
---------